OpenStudy (anonymous):

MEDAL AND FAN Write the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3 x - 7.

OpenStudy (gorv):

perpendicular means product of slope of two line =-1 1/3 *slope2=-1 slope2=-3

OpenStudy (gorv):

y-(-2)=-3(x-4) y+2=-3x+12 y=-3x+10

OpenStudy (gorv):

@Kstate

OpenStudy (anonymous):

It is y=-3x+10

OpenStudy (gorv):

yep thats wat i ...

OpenStudy (anonymous):

got

OpenStudy (gorv):

:)